# A Notion of Complexity for Theory of Mind via Discrete World Models

### v2

@@ 3.2.2 On the Complexity of DWM
DWM progressively calls an LLM $T>0$ times to generate informative states. For a ToM problem of length $n$ (i.e., the number of input tokens), which we assume, w.l.o.g., that can be split into $k$ chunks of approximately the same length $|x\oplus p_{i}\oplus w|=\frac{n}{T}$, the number of tokens generated by an LLM is in the order of $\mathcal{O}(\sum_{t=1}^{T}|x\oplus p_{t}\oplus w\oplus a_{t}|^{2})$, where $p_{t}$ ($a_{t}$) is the portion of the problem (answer) prompted (retrieved) at iteration $t$. With the further assumption that each answer retrieved at split $t\leq T$ has the same length $o$, the complexity is further simplified to be asymptotic to $\mathcal{O}((\frac{n}{T}+o)^{2})$. Compared to CoT, whose complexity is $\mathcal{O}(n+o)$, DWM requires an additional linear number of calls to the model. On the other hand, ToT with the same number of splits $\frac{n}{T}$ and $m>1$ experts results in even higher complexity, i.e., asymptotic to $\mathcal{O}(m(\frac{n}{T}+o)^{2})$.
@@ 4 Experimental Evaluation
The experiments are organised as follows. We first test the performance of DWM on ToMi Le et al. (2019), MindGames Sileo and Lernould (2023), Adv-CSFB  Shapira et al. (2023), SocialIQA Sap et al. (2019), and FANToM Kim et al. (2023), comparing it with CoT Wei et al. (2022), ToT Yao et al. (2023) and prompting with structured data (struct), i.e., the model is queried to first represent the problem in a structured format such as JSON or Yaml. We further show that ToMi has been memorised word for word by GPT models, with CoT (and any technique that leaves the input unchanged) being the best-performing method. We then quantify the complexity of the benchmarks introduced above and highlight the correlation with the models’ performances. Our framework shows complexity ranges between easy and hard problems, even within a benchmark. We conduct our experiments on GPT-3.5-Turbo, GPT-4 OpenAI (2023), LLaMA3-70B AI@Meta (2024) and Mixtral 8x7B Jiang et al. (2024).
@@ 4.1 DWM on ToM Benchmarks
We report results for GPT-3.5-Turbo and Mixtral 8x7B on the five ToM benchmarks: for reasons of space, results for LLaMA3-8B, LLaMA3-70B and GPT-4 are reported in the Appendix, Section B.1. As illustrated in Figure 4 (top), DWM improves the performance of GPT-3.5-Turbo on Mindgames, FANToM and Adv-CSFB by a solid margin. On SocialIQa, which has very short inputs, DWM performs slightly worse than CoT but better than ToT. On the other hand, on ToMi, the best prompting techniques are CoT and ToT. As analysed in the next section, we believe this not be caused by memorisation, which we prove happening, but by the inherent complexity of the task. With Mixtral 8x7B (Fig. 4 (bottom)), DWM improves the performance on ADVcsfb, FANToM, ToMi and Mindgames, and pairs that of CoT on SocialIQa.
@@
We qualitatively analysed the information elicited by an LLM when prompted with DWM and discovered that it forces a model to output information not explicitly available in the prompt. Consider the ToMi example in Figure 5 where GPT-4 is prompted with a situation where agents interact and are then queried with the first-order belief of Benjamin. With CoT, the model makes an erroneous assumption about the presence of Benjamin and Isabella in the room. On the other hand, when prompted with DWM, GPT-4 provides an informative description of each state space, particularly the knowledge and the uncertainty of each agent’s beliefs, and eventually answers correctly. One example per benchmark is available in the Appendix, Section B.2, while many more are available for inspection in the Code Supplementary Material. Such phenomenon is ubiquitous to all the ToM tasks we tested, a hint that DWM elicits the ToM capabilities of LLMs without requiring external information or solvers.
@@
Recent works expressed concern about ToM benchmarks’ efficacy in memorisation Jacovi et al. (2023); La Malfa et al. (2024). This motivated us to quantify and then analyse the impact of memorisation of ToM benchmarks on performance. We computed the percentage of memorised prompts to understand whether that affects the performance of techniques, such as DWM, that split the prompt into chunks and introduce additional information instead of CoT, which leaves the input prompt unchanged. As illustrated in Table 1, ToMi and FANToM have been heavily memorised, with entire portions of the benchmarks that can be retrieved word for word from GPT-3.5-Instruct (the autocomplete model by OpenAI). Despite that, no clear evidence of a performance drop in DWM induced by memorisation exists. For GPT-3.5, despite CoT having higher performance on ToMi, DWM is better on FANToM (Figure 4). We hypothesise that as long as a memorised problem is prompted, either in its exact form (as for CoT) or split as in DWM, the most potent models can recover it alongside the ground truth label, thus invalidating the test for both. We conclude with a note of caution. While we discovered that ToMi and FANToM are memorised by GPT-3.5-Instruct, that doesn’t imply any LLM, including GPT-3.5-Turbo and GPT-4, whose training details are not released publicly, has been trained on that data.
@@ 5 Conclusions
Training and evaluating on the same dataset produce positively biased data on the model’s performance. While running our benchmarks on ToMi, we discovered that the GPT-3.5 model had completely memorized parts of the dataset. This motivated us to extend the memorization test to the other tasks. We urge the research community to include a memorization section on every benchmark study with public datasets used in their works. This data is crucial to conduct fair and unbiased research on evaluating LLMs’ abilities Jacovi et al. (2023). Future works will include an analysis of the memorisation rate of other ToM tasks alongside tests to quantify their impact on different models.
@@
The datasets and pre-trained LLMs that we use are all publicly available. This paper focuses on ToM problems’ hardness and prompting methods. We highlight that LLMs do not guarantee the production of factual data or correct reasoning steps, and the prompting methods developed here should not be regarded as the source of truth in making decisions.
Acknowledgments
ELM is supported by the Alan Turing Institute. AGC is supported by the Economic and Social Research Council (ESRC) under grant ES/W003473/1, by the Fundamental Research priority area of The Alan Turing Institute, and by the Turing Defence and Security programme through a partnership with the UK government in accordance with the framework agreement between GCHQ and the Alan Turing Institute. The support of Microsoft under their Accelerating Foundation Models Research initiative in providing Azure credits to AGC is gratefully acknowledged.
@@ A.1 Experimental Details
Most of the language models used in this work follow the Language Models as a Service (LMaaS) paradigm La Malfa et al. (2023). This model of service does not allow transparency and hinders reproducibility. Reproducibility is difficult to achieve as common software development frameworks, such as CI/CD pipeline, ease the update of the public service, but change the underlying entity. From this follows that the model tested by the researcher could change at any time. This is not solvable from the outside. Researchers have no control over the software engineering practices inside a LMaaS, but could set some parameters to offer the highest possible grade of reproducibility. We set the temperature to zero or enable greedy decoding by default (this does not imply determinism even if model weights are not changed).  [^1] The main explanation are the https://github.com/pytorch/pytorch/issues/75240”non deterministic cuda cores” another could ”be batched inference in sparse MoE models”, see https://152334h.github.io/blog/non-determinism-in-gpt-4/here In prompting methods where the creativity of the response is exploited for better performance, e.g., Tree of Thoughts Yao et al. (2023), we set the temperature to $0.7$, the value proposed in the reference papers.
@@
We use Huggingface for Mixtral 8x7B. Groq Cloud for LLama-3-7B and LLama-3-70B. Microsoft sponsorship for GPT-3.5 and GPT-4 access.
@@ B.2 DWM Elicits More Informed Mental States in LLMs
Figure 9: Example of a real ToMi example where GPT-4 fails when prompted with CoT, yet succeeds with DWM. CoT elicits an untruthful reasoning process (in red), while DWM correctly informs the model with the correct information about Benjamin’s first-order belief (in green).

### v3

@@ 3.2.2 On the Complexity of DWM
DWM progressively calls an LLM $T>0$ times to generate informative states. For a ToM problem of length $n$ (i.e., the number of input tokens), which we assume, w.l.o.g., that can be split into $k$ chunks of approximately the same length $\lvert x\oplus p_{i}\oplus w\rvert=\frac{n}{T}$, the number of tokens generated by an LLM is of the order of $\mathcal{O}(\sum_{t=1}^{T}\lvert x\oplus\left(\bigoplus_{i=1}^{t-1}p_{i}\oplus w\oplus a_{i}\right)\rvert)$, where $p_{t}$ ($a_{t}$) is the portion of the problem (answer) prompted (retrieved) at iteration $t$. With the further assumption that each answer retrieved at split $t\leq T$ has the same length $o$, the complexity is further simplified to be asymptotic to $\mathcal{O}((\frac{n}{T}+o)^{2})$. Compared to CoT, whose complexity is $\mathcal{O}(n+o)$, DWM requires an additional linear number of calls to the model. On the other hand, ToT with the same number of splits $\frac{n}{T}$ and $m>1$ experts results in even higher complexity, i.e., asymptotic to $\mathcal{O}(m(\frac{n}{T}+o)^{2})$.
@@ 4 Experimental Evaluation
The experiments are organised as follows. We first test the performance of DWM on ToMi Le et al. (2019), MindGames Sileo and Lernould (2023), Adv-CSFB  Shapira et al. (2024), SocialIQA Sap et al. (2019), and FANToM Kim et al. (2023), comparing it with CoT Wei et al. (2022), ToT Yao et al. (2024) and prompting with structured data (struct), i.e., the model is queried to first represent the problem in a structured format such as JSON or Yaml. We further show that ToMi has been memorised word for word by GPT models, with CoT (and any technique that leaves the input unchanged) being the best-performing method. We then quantify the complexity of the benchmarks introduced above and highlight the correlation with the models’ performances. Our framework shows complexity ranges between easy and hard problems, even within a benchmark. We conduct our experiments on GPT-3.5-Turbo, GPT-4 OpenAI (2023), LLaMA3-70B AI@Meta (2024); Dubey et al. (2024) and Mixtral 8x7B Jiang et al. (2024).
@@ 4.1 DWM on ToM Benchmarks
We report results for GPT-3.5-Turbo and Mixtral 8x7B on the five ToM benchmarks: for reasons of space, results for LLaMA3-8B, LLaMA3-70B and GPT-4 are reported in the Appendix, Section B.1. As illustrated in Figure 4 (top), DWM improves the performance of GPT-3.5-Turbo on Mindgames, FANToM and Adv-CSFB by a solid margin. On SocialIQa, which has very short inputs, DWM performs slightly worse than CoT but better than ToT. On the other hand, on ToMi, the best prompting techniques are CoT and ToT. While one may think memorisation plays a role in boosting the performance of LLMs with these prompting techniques, in the next section, we provide evidence this hypothesis is not necessarily true. With Mixtral 8x7B (Fig. 4 (bottom)), DWM improves the performance on ADVcsfb, FANToM, ToMi and Mindgames, and reaches that of CoT on SocialIQa.
@@
We qualitatively analysed the information elicited by an LLM when prompted with DWM and discovered that it forces a model to output information not explicitly available in the prompt. Consider the ToMi example in Figure 5 where GPT-4 is prompted with a situation where agents interact and are then queried with the first-order belief of Benjamin. With CoT, the model makes an erroneous assumption about the presence of Benjamin and Isabella in the room. On the other hand, when prompted with DWM, GPT-4 provides an informative description of each state space, particularly the knowledge and the uncertainty of each agent’s beliefs, and eventually answers correctly. One example per benchmark is available in the Appendix, Section B.2, while many more are available for inspection in the Code Supplementary Material. This phenomenon is ubiquitous to all the ToM tasks we tested, a hint that DWM may elicit the ToM capabilities of LLMs without requiring external information or solvers.
@@
Recent works expressed concern about ToM benchmarks’ efficacy in memorisation Jacovi et al. (2023); La Malfa et al. (2024b). This motivated us to quantify and then analyse the impact of memorisation of ToM benchmarks on performance. We computed the percentage of memorised prompts to understand whether that affects the performance of techniques, such as DWM, that split the prompt into chunks and introduce additional information instead of CoT, which leaves the input prompt unchanged. As illustrated in Table 1, ToMi and FANToM have been heavily memorised, with entire portions of the benchmarks that can be retrieved word for word from GPT-3.5-Instruct (the autocomplete model by OpenAI). Despite that, no clear evidence of a performance drop in DWM induced by memorisation exists. For GPT-3.5, despite CoT having higher performance on ToMi, DWM is better on FANToM (Figure 4). We hypothesise that as long as a memorised problem is prompted, either in its exact form (as for CoT) or split as in DWM, the most potent models can recover it alongside the ground truth label, thus invalidating the test for both. We conclude with a note of caution: while we discovered that ToMi and FANToM are memorised by GPT-3.5-Instruct, that does not imply any LLM, including GPT-3.5-Turbo and GPT-4, whose training details are not released publicly, has been trained on that data.
@@ 5 Conclusions
Training and evaluating on the same dataset produces positively biased data on the model’s performance. While running our benchmarks on ToMi, we discovered that the GPT-3.5 model had completely memorized parts of the dataset. This motivated us to extend the memorization test to the other tasks. We urge the research community to include a memorization section on every benchmark study with public datasets used in their works. This data is crucial to conduct fair and unbiased research on evaluating LLMs’ abilities Jacovi et al. (2023). Future works will include an analysis of the memorisation rate of other ToM tasks alongside tests to quantify their impact on different models.
On element interactivity.
Sweller Sweller (2010) proposes a measure of complexity for cognitive tasks that encompasses three main components, namely the intrinsic, extraneous, and germane cognitive load. In its framework, which has wide applications in education, the intrinsic load relates to the number of references, or interactions, between the elements of a problem, i.e., the information or concept that needs to be understood to answer the question. Our framework approximates the intrinsic and extraneous loads to be single sentences in a ToM problem, which is not assured to be the best measure.
@@
The datasets and pre-trained LLMs that we use are all publicly available. This paper focuses on ToM problems’ hardness and prompting methods. We highlight that LLMs do not guarantee the production of factual data or correct reasoning steps, and the prompting methods developed here should not be regarded as the source of truth in decision-making.
@@ A.1 Experimental Details
Most of the language models used in this work follow the Language Models as a Service (LMaaS) paradigm La Malfa et al. (2024a). This model of service does not allow transparency and hinders reproducibility. Reproducibility is difficult to achieve as common software development frameworks, such as CI/CD pipeline, ease the update of the public service but change the underlying entity. From this, it follows that the model tested by the researcher could change at any time. This is not solvable from the outside. Researchers have no control over the software engineering practices inside a LMaaS, but could set some parameters to offer the highest possible grade of reproducibility. We set the temperature to zero or enable greedy decoding by default (this does not imply determinism even if model weights are not changed).  [^2] The main explanation is the https://github.com/pytorch/pytorch/issues/75240”non-deterministic cuda cores” another could ”be batched inference in sparse MoE models”, see https://152334h.github.io/blog/non-determinism-in-gpt-4/here In prompting methods where the creativity of the response is exploited for better performance, e.g., Tree of Thoughts Yao et al. (2024), we set the temperature to $0.7$, the value proposed in the reference papers.
@@
We use Huggingface for Mixtral 8x7B. Groq Cloud for LLama-3-8B and LLama-3-70B. Microsoft sponsorship for GPT-3.5 and GPT-4 access.
@@ B.2 DWM Elicits More Informed Mental States in LLMs
Figure 9: Example of a a ToMI instance where GPT-4 fails when prompted with CoT, yet succeeds with DWM. CoT elicits an untruthful reasoning process (in red), while DWM correctly informs the model with the correct information about Benjamin’s first-order belief (in green).